T = (((X * A * B * (100 - C) / 100) - Y) / Z) ... solve for X

Can you help me with the answer? Thank you.

Answer by jim_thompson5910(35256)

T = (((X * A * B * (100 - C) / 100) - Y) / Z)

ZT = (X * A * B * (100 - C) / 100) - Y

ZT + Y = (X * A * B * (100 - C) / 100)

100*(ZT + Y) = X * A * B * (100 - C)

100ZT + 100Y = X * A * B * (100 - C)

(100ZT + 100Y)/(A * B * (100 - C)) = X

X = (100ZT + 100Y)/(A * B * (100 - C))